### v3

# LongGenbench: Benchmarking Long-Form Generation in long-context LLMs

Abstract
The abilities of long-context language models (LMs) are often evaluated using the ”Needle-in-a-Haystack” (NIAH) test, which comprises tasks designed to assess a model’s ability to identify specific information (”needle”) within large text sequences (”haystack”). While these benchmarks measure how well models understand long-context input sequences, they do not effectively gauge the quality of long-form text generation—a critical aspect for applications such as design proposals and creative writing. To address this gap, we have introduced a new long-form text evaluation benchmark, LongGenBench, which tests models’ ability to identify specific events within generated long text sequences. In this benchmark, we prompt long-context LMs to create long-form text that must include particular events or constraints and evaluate their ability to incorporate these elements. We evaluated ten long-context LMs across four distinct scenarios, three types of prompt instructions, and two different generation-length settings (16K and 32K). Although these models perform well on NIAH benchmarks, none demonstrated satisfactory performance on the LongGenBench , raising concerns about their ability to generate coherent long-form text that follows instructions. Additionally, as the length of the generated text increases, all models exhibit a significant drop in performance [^1] We have open-sourced LongGenBench to support the comprehensive evaluation of long-context LMs in long-form text generation. Our code and data are available at https://github.com/mozhu621/LongGenBench/..
1 Introduction
Recent advances in large language models (LLMs) have significantly improved their ability to process and generate long text sequences, which is essential for various natural language processing tasks (Xiong et al., 2023; Bai et al., 2023a). For instance, GPT-4  (Achiam et al., 2023) manages context windows of up to 128K tokens, Claude 2.1 (Anthropic, 2024a) supports up to 200K tokens, Claude 3 series (Anthropic, 2024b) handles inputs exceeding 1 million tokens, and Gemini 1.5 (Reid et al., 2024) processes context windows that span several million tokens. As LLMs become increasingly capable of processing long input text sequences, various methods have been proposed to evaluate their effectiveness in handling extended contexts. Existing benchmarks employ synthetic or hybrid evaluation, such as passkey retrieval (Mohtashami & Jaggi, 2023), needle-in-a-haystack (Kamradt, 2023), Longbench (Bai et al., 2023a), Ruler (Hsieh et al., 2024), and Needlebench (Li et al., 2024), to evaluate the retrieval, understanding, and reasoning capabilities of LLMs. However, currently there are no automated benchmarks for evaluating long output text sequences, which overlooks the quality aspect in long-form text generation tasks.
Generating long-form text sequences is crucial for various natural language processing tasks, such as document summarization(Kumar et al., 2024), creative writing(Hua & Wang, 2020; Hu et al., 2022) and comprehensive question answering(Stelmakh et al., 2022; Lee et al., 2023; Bai et al., 2024). However, evaluating these long sequences poses significant challenges. First, there is a lack of evaluation datasets, as most existing ones are designed for shorter text output ($\leq$ 2K tokens) (Fan et al., 2018; 2019a; Dasigi et al., 2021), making them less suitable for super-long-form generation  ($\geq$ 16K tokens) (Bai et al., 2024). Furthermore, as the input context window gets increasingly longer, long-form generation tasks will naturally produce even longer sequences, which amplifies the gap found in the existing dataset. Second, there is no robust method for reliably assessing the quality of these long text sequences. Researchers often rely on manual evaluations, which, while thorough, are costly and impractical for large-scale assessments. Alternatively, researchers resort to the approach “LLM-as-a-judge” (Zheng et al., 2024), which uses models capable of processing long inputs to evaluate text quality. However, the scores generated by these black-box models can be difficult to explain and may not align with human expectations, raising concerns about their validity. These challenges highlight the need for a more specialized dataset with reliable evaluation methods for super-long-form text generation.
This study introduces LongGenBench , a new benchmark assessing super-long-form generation capabilities in a long-context language model. This benchmark evaluates the quality of generated content and adherence to instructions, demanding varying levels of compliance based on the specific instructions. For instance, creative writing tasks like storytelling or diary entries for fictional characters might involve detailing significant events such as family birthdays or annual vacations. Meanwhile, project planning tasks, such as urban planning, could require the inclusion of essential constraints in the document. To establish a robust benchmark, we have designed four distinct scenarios: Diary Writing, Menu Design, Skyscraper Design, and Urban Planning. Each scenario comprises three sub-scenarios with specific task instructions—single instance, range, and periodicity. These sub-scenarios capture and reflect the complexity behind long-form generation tasks, where long sequence texts must include and adhere to various criteria. These scenarios thus serve as a foundational base for evaluating long-form generation.
Our benchmark provides a fine-grained evaluation of generated text sequences by assessing for the presence of essential events or constraints specified in the prompt instructions. Unlike the NIAH (Kamradt, 2023), which tests the fundamental capability of long-context LMs to process lengthy input sequences and identify critical information, our benchmark focuses on instruction-following in super-long-form generation tasks. Specifically, it challenges models to include specific events or constraints, akin to verifying the needle in the output text sequence. This method not only captures and reflects the overall quality of the task but also minimizes the need for human evaluation or excessive reliance on LLM-based judgments. Our study findings indicate that current long-context language models are inadequate when handling LongGenBench, even with texts shorter than 16K (see Figure 1).
Figure 1: The left side of the figure displays the average performance results of the model at Ruler (16K) (Hsieh et al., 2024), whereas the right side shows the model’s average performance on the LongGenBench  (16K).
In summary, our major contributions are as follows:
  • To the best of our knowledge, this is the first study to introduce the challenge of super-long-form generationin long-context language models. The generation, retrieval, comprehension, and reasoning within these extended contexts are crucial.
  • We introduce LongGenBench, a dataset that includes tasks for evaluating the super-long-form generation  capabilities of LLMs across multiple text token intervals (16K and 32K) and various text depth ranges.
  • We conducted comprehensive experiments on both open-source and close-source models. Our findings indicate that most models fail to complete even a fraction of the super-long-form generation tasks they are purported to manage.
2 LongGenBenchBenchmark
In this section, we provide a detailed examination of our benchmark as outlined, encompassing Task definition, Different Scenario setups, Specific task instruction, Evaluation metric and Evaluations pipeline.
Figure 2: LongGenBench Overview: 1) Scenario Selection: Select from four scenarios—Diary, Menu Design, Skyscraper Design, and Urban Planning—each offered in both short and long versions to determine the main task prompt. 2) Task Instruction: Employ the template libraries SI (Single), RI (Range), and PI (Periodic) to generate tasks characterized by random times or locations, along with the corresponding prompts and verification sets. 3) Instruction Synthesis: Integrate all prompts generated in the prior step to create a comprehensive set of instructions with a final check-set. 4) Example: An illustration of Sophia’s weekly diary task is provided as an example.
2.1 Task Definition
Evaluating the quality of super-long-form generation  is both crucial and challenging, given the inherent complexity of long texts. Traditional human evaluation methods, while precise, are expensive and not scalable. Although feasible, using large language models for assessment often lacks interpretability. Thus, we focus on the ”instruction following” task in super-long-form generation . In this study, We define a task type termed Strictly Sequential Tasks, which involves the sequential completion of subtasks $\mathbf{T}=\{T_{1},T_{2},T_{3},\ldots,T_{n}\}$, where each subtask is responsible for generating a specific volume of text. For instance, an LLM might be tasked with designing a 100-floor skyscraper, specifying the content and purpose of each floor.
2.2 Four Distinct Scenario Setups
In conjunction with the design of a 100-floor skyscraper, we have devised three distinct task scenarios to supplement our predefined tasks, as illustrated in Figure 2 (1). These scenarios fall into two categories: Temporal (Diary Writing, Menu Design) and Spatial (Skyscraper Design, Urban Planning). Moreover, each scenario incorporates both short and long version to assess the effectiveness of various output lengths. Each scenario is carefully engineered to evaluate the model’s capacity to produce context-sensitive and detailed lengthy texts tailored to our specific task requirements. Table 1 offers comprehensive descriptions for each scenario, with each designed around a unique template to generate 100 different task instructions [^2] Examples of Task Instructions for each scenario are provided in Appendix B..
Table 1: Detailed Scenario Descriptions
Category Scenarios Task Task Description
Temporal Diary Weekly Diary Generate entries for each week of the year
Daily Diary Generate entries for each day of the year
Menu Weekly Menu Plan menus for each week of the year
Daily Menu Plan menus for each day of the year
Spatial Skyscraper Design 100-floor Design Develop a design for a 100-floor skyscraper
361-floor Design Develop a design for a 300-floor skyscraper
Urban Planning 10x10 block Design Design an urban layout on a 10x10 block grid
19x19 block Design Design an urban layout on a 19x19 block grid
2.3 Specific Task Instruction
To enhance task control and flexibility, we have developed three distinct task settings:
  • Single Instruction (SI): Introduces specific information uniquely within the text.
    $T_{S}=\{T_{s_{1}},T_{s_{2}},\ldots\}$
  • Range Instruction (RI): Incorporates specific information at designated ranges within the text.
    $T_{R}=\{T_{i},T_{i+1},\ldots,T_{i+j}\}$
  • Periodic Instruction (PI): Periodically disperses information at predefined intervals throughout the text.
    $T_{P}=\{T_{n},T_{2n},\ldots,T_{m*n}\}$
  • Check Set: Includes tasks for all three aforementioned settings.
    $check\_set=\{T_{S},T_{R},T_{P}\}$
For instance, in the design of a 100-floor skyscraper, the Single Instruction may specify that the 34th floor hosts an aerial gym and the 54th floor houses a law firm. The Range Instruction might designate floors 1 through 9 as a comprehensive shopping mall, whereas the Periodic Instruction could dictate that starting from the 20th floor, every 10th floor incorporates a small aerial garden. We utilize over 20 templates for each type of instruction, with the floors or locations being randomly assigned to ensure task diversity. These settings, applied via various templates, guarantee controlled coverage across all textual positions, thus facilitating a comprehensive and efficient evaluation as illustrated in Figure 2 (2). For instance, my wife’s birthday falls on October 8th, and our family trip to Miami is scheduled from December 24th to celebrate Christmas, among other events. Through this approach, we generate the main task instructions $T$ and simultaneously acquire the corresponding $check\_set$, which supports subsequent evaluations and constructs a task conducive to super-long-form generation. Subsequently, we splice the main task prompt with the specific task instructions (STI) [^3] Each of our main task instructions $T$ splice 5 single instructions, 1 range instruction task, and 1 periodic instruction task. and add the generation prompt to form the final evalution data.
2.4 evaluation Metric
To quantitatively evaluate performance for LongGenBench tasks, we introduce three complementary metrics:
Main Task Completion
This metric evaluates the extent to which all designated subtasks are accomplished. The completion rate is quantified using the following equation:
Completion Rate=Number of Completed SubtasksTotal Number of $\text{Completion Rate}=\frac{\text{Number of Completed Subtasks}}{\text{Total Number of Subtasks}}\times 100\%$
In this context, the numerator denotes the count of subtasks successfully executed by the model, and the denominator represents the total number of subtasks defined in the Strictly Sequential Task. For instance, does the model consistently complete a diary entry for each day without omitting any dates?
Specific Task Instruction Completion
To assess the performance of specific subtasks under defined conditions, we employed the Llama3.1-8B instruction model (Dubey et al., 2024) for a few-shot evaluation. This method allowed us to verify compliance of each subtask with its corresponding Specific Task Instructions, effectively transforming the assessment into a binary classification task (Completed” and ”Uncompleted). Additionally, we manually annotated 500 data points to validate the consistency of the outcomes from the Llama3.1-8B-instruct model with human assessments [^4] Detailed evaluation prompts can be found in the Appendix C.. . We then calculated the completion counts for the Single Instruction (SI), Range Instruction (RI), and Periodic Instruction (PI). To assess the performance of the SPI (Specific Task Instruction) process, we computed the ratio of the total number of relevant sub-scenarios included in the answers to the number of completed instructions, denoted as STIC-1:
[…]
Additionally, we calculated the total number of sub-scenarios across all Specific Task Instructions. This enumeration supported the calculation of ratios for SI, RI, PI, to the total Specific Task Instructions, resulting in the Specific Task Instruction Completion-2 metric (STIC-2):
[…]
STIC-1 is primarily concerned with the completion rate of instructions that result in sub-scenarios, focusing on whether instructions are completed. In contrast, STIC-2 assesses the overall completion of the SPI task, including the presence of sub-scenarios and their completion status.
2.5 evaluations Pipeline
Our evaluation process consists of the following steps: 1) First, we use a long-context language model (LLM) to complete the aforementioned task instruction $T$ and obtain its Answer $A$, which is then divided into sub-tasks as $A=\{A_{1},A_{2},\ldots,A_{n}\}$. 2) Next, based on the instructions that need to be evaluated in the check_set, we identify the corresponding sub-tasks in the Answer. 3) Finally, we evaluate each sub-task by $\text{eval}(A_{i},T_{i})$ to calculate the final completion score, as specified by the algorithm 1.
Algorithm 1 evaluations Pipeline
1:Initialization:
2:Task $\text{Task instructions}\rightarrow T$
[…]
5:
6:Main Process:
7:Use Tested model to get Answer for $\text{Use Tested model to get Answer for }T\rightarrow A$
8:$A\rightarrow\{A_{1},A_{2},\ldots,A_{m}\}\text{, split into subtasks}$
9:empty set for storing $\text{empty set for storing evaluations}\rightarrow E$
10:for each $T_{i}$ in $Check\_Set$ do
11:     if there is $A_{i}\text{ matching }T_{i}$ then
12:         $\text{eval}(A_{i},T_{i})\rightarrow E_{i}$
13:         $E\rightarrow\text{Add }E_{i}\text{ to }E$
14:     end if
15:end for
16:$\sum E\rightarrow Score\text{, compute the final completion score}$
17:return $Score$
3 Experiment
3.1 Experimental Setup
Models & Inference setup
We selected ten long-context LLMs, comprising eight open-source models and two closed-source model, covering a range of model sizes from 7B to 72B with Mixture of Experts (MoE) architecture and claimed context lengths ranging from 32k to 128K tokens. Detailed specifications of these models are provided in Appendix A. We evaluated all models using the vLLM (Kwon et al., 2023) serving system, which features efficient KV cache memory management. Inference was conducted with BFloat16 precision on eight NVIDIA A800 GPUs, using greedy decoding.
Task Configurations
For each scenario, we generated 800 examples for the specified lengths (16k, 32k) according to the required chat template for each model. To ensure the models produced relevant responses without generating off-topic content or refusing to answer, we prefixed each task input with a specific answer prompt and verified the presence of the target output using recall-based accuracy measures.
3.2 Main Result
Models Claimed Short-version (16K) Long-version (32K)
Length CR STIC-1 STIC-2 Len. wAvg CR STIC-1 STIC-2 Len. wAvg
Models with 7-10B Parameters
Mistrial-v0.2-7B 32K 81.8% 22.0% 17.4% 7296 14.3% 48.2% 37.0% 16.3% 16146 7.9%
LLama3.1-8B 128K 93.5% 23.4% 22.0% 8804 20.6% 77.6% 26.5% 18.9% 17354 14.6%
Qwen2-7B 128K 60.0% 27.9% 16.1% 5138 9.7% 40.0% 31.7% 12.6% 9617 5.0%
FILM-7B 128K 36.0% 9.9% 3.9% 6280 1.4% 37.4% 30.9% 10.9% 13775 4.1%
LongWriter-llama3.1-8B 128K 46.0% 22.6% 9.8% 11036 4.5% 34.5% 33.6% 10.0% 19925 3.5%
Models Larger Than 20B Parameters
Mixtral-8x7B 32K 83.0% 35.4% 28.0% 8113 23.3% 60.5% 39.9% 22.3% 15839 13.5%
LLama3.1-70B 128K 79.3% 24.9% 21.1% 8055 16.7% 63.1% 35.8% 21.7% 15197 13.7%
Qwen2-72B 128K 94.3% 25.5% 24.0% 8013 22.7% 66.2% 27.5% 17.4% 19845 11.5%
Closed-source Model
GPT-4o-mini 128K 97.0% 29.0% 27.9% 8940 26.9%
GPT-4o 128K 67.2% 34.9% 19.9% 9055 12.5%
Table 2: Long-form generation Performance (%) of selected models evaluated at length from 16k and 32k. The weighted average score (wAvg) is the product of CR and STBC-2, used to represent the model’s final performance at the given task length. Note that the GPT-4-32K is currently closed for use, and the longest versions that can be used are the GPT-4o and GPT-4o-mini 16K output limitation.
Completion Rate:
Significant disparities in performance across models primarily stem from differences in architecture and training datasets. Notably, models with varying parameter sizes, such as Llama3.1-8B-instruction (Dubey et al., 2024) (under 10 billion parameters), Qwen-72B (Yang et al., 2024) (over 20 billion parameters), and GPT-4o-mini (OpenAI, 2024a) (a closed-source model), have demonstrated superior efficacy, successfully completing most primary tasks in full. In contrast, some models struggle with these tasks, exhibiting limitations such as: 1) models responding solely to specified subtasks, neglecting others, and 2) models halting after only completing the initial task segment, despite prompts requiring full sequential subtask completion. This issue may originate from the current instructional tuning data, which could cause partial responses in complex, lengthy tasks. Especially in GPT-4o(OpenAI, 2023), it recognizes that this task will generate a long output and only provides a few examples.
STIC-1 & STIC-2:
The STIC-1 metric demonstrates the robust performance of both the Mixtral-8*7B (Mistral.AI, 2023) and GPT-4o (OpenAI, 2024a) models in adhering to the Scenario Task Instruction (STI) during sub-scenario queries. However, an analysis of the overall STI completion rates reveals subtle performance discrepancies. Models with lower completion rates experience a marked decline from STIC-1 to STIC-2. In this metric, the Mixture of Experts (MOE) models outperformed their dense counterparts.
Length (Number of words):
We calculate the average output word count for models that consistently complete all subtasks, achieving at least an 80% completion rate in sub-scenarios, excluding data from unsuccessful attempts. Most models substantially exceed previous benchmarks for long-form generation tasks in terms of output length. Notably, the LongWriter (Bai et al., 2024) model excels, efficiently meeting word count requirements for each subtask.
Given the results and the weighted average score (wAvg) at a sequence length of 16K, the open-source Qwen2-72B and the closed-source GPT-4o models demonstrate optimal performance. At a sequence length of 32K, the LLama3.1-8B model, known for its excellent completion rate, outperforms models with larger parameters, highlighting its efficiency in managing extended lengths.
3.3 Accuracy Trend with Varying Sequence Length
As show in Fig. 3, there is a noticeable decline in model performance over time. As the outputs become longer, all models gradually fail to adhere to the initial instructions, resulting in a performance degradation analogous to trends observed in the NIAH dataset. This deviation is particularly pronounced when the output exceeds 4,000 tokens, where adherence to instructions significantly diminishes, with marked deterioration at the 4,000-token threshold and substantial ineffectiveness by 16,000 tokens. Conversely, in tasks such as those involving NIAH and more complex multi-needle tasks of similar lengths, these models demonstrate near-perfect performance. Such disparities in performance cannot be easily explained by prevailing theories, such as ”loss in the middle” (Liu et al., 2024a), especially since the instructions are positioned within the first 500 tokens.
Figure 3: The right side of the figure illustrates the model’s performance on STI tasks at 16K as sequence length increases, whereas the left side depicts performance at 32K. All curves have been smoothed with a Moving Average.
3.4 Three Specific Task Instruction
In Figure 4, we present the performance metrics across various tasks. The model shows similar, yet not identical, proficiency in single and range tasks, suggesting its effective execution of direct instructions for these task types. This reduction in performance can be ascribed to the additional cognitive load associated with interpreting periodic instructions, such as identifying relevant weeks in statements like ”every four weeks starting from week 10.” Consequently, outcomes for periodic tasks are substantially poorer compared to those for single and range tasks, which have well-defined parameters. Typically, the model’s performance hierarchy is represented as $single>range>periodic$.
3.5 Compare with long context input
In this subsection, we examine the relationship between a model’s ability to handle long-range inputs and its performance on long-range outputs. Specifically, we investigate whether a model’s capacity to manage long-range inputs corresponds to improved performance on long-range outputs. For this analysis, we use the RULER dataset. RULER is a novel synthetic benchmark designed to flexibly configure sequence length and task complexity, making it ideal for comprehensive evaluations of long-context language models. We compare the models’ performance on sequences of the same length, as shown in Figure 4, which indicates a certain degree of correlation between input handling and output performance. At 16K, the Pearson correlation coefficient is 0.51, while at 32K, it increases to 0.66, highlighting a significant overlap in their capabilities.
(a) Performance Comparison of Three STI Type
(b) Performance Comparison on Ruler and LongGenBenchTasks
Figure 4: The left fig displays the performance of three different STI models, with the red line representing the average for each category. The right fig shows the performance and correlation of the ruler and LongGenBench  at the same length settings.
4 analysis and limitaion
Richness of Content
Even when sub-scenarios are specifically designed to enhance task richness, the model’s outputs tend to converge as output volume increases. This results in a homogeneity of recorded events despite differences in time and location. Such convergence not only degrades performance, but also diminishes output diversity. Attempts to mitigate this issue by adjusting the parameter $repetition\_penalty$ during model inference have not been effective in our experiments.
Rationality of Content
Our research presently focuses only on evaluating instruction-following capabilities within tasks; a more comprehensive analysis of task rationality and integrity could be a direction for future research. For instance, if a virtual task pertains to a specific profession, all job-related activities recorded in a diary should correspond to that profession. Additionally, we noted that temperature records in diaries fail to reflect realistic temporal changes. For example, in San Francisco, one would expect temperatures to rise from cold (0-10 degrees Celsius) at the beginning of the year to warmer (20-30 degrees Celsius) by mid-year; however, our findings indicate temperatures remain in the warmer range (20-30 degrees Celsius) even into December.
Instruction Data
Significant discrepancies in the performance of most models compared to their ability to process inputs, such as Ruler(Hsieh et al., 2024), are likely due to the length distribution of instruction tuning data. These data sets are typically brief, usually under 200 tokens, and lack extended instructional content. Organizing or synthesizing instruction tuning data with longer outputs could be a potential solution for future research endeavors.
5 Related Work
Long-context Benchmarks and Tasks.
Our work is closely related to other works on benchmarking long-context language models. ZeroSCROLLS (Shaham et al., 2023) covers ten realistic natural language tasks, such as long-document QA and (query-based) summarization. L-Eval (An et al., 2024a) also uses realistic data, which was filtered manually to ensure quality. LongBench (Bai et al., 2023b) contains tasks in a bilingual setting. InfiniteBench (Zhang et al., 2024) includes tasks with length greater than 100K tokens. LTM (Castillo et al., 2024) targets the evaluation of long-term conversations. To isolate the effect of parametric knowledge, previous works (Dong et al., 2023; Li et al., 2023b) also propose to use documents posted online later than a certain cutoff date, or leverage extremely low-resource materials (Tanzer et al., 2024). Compared to realistic benchmarks, synthetic tasks are more flexible to control the setup (e.g., sequence length and task complexity) and less affected by parametric knowledge. Recent works have mostly focused on retrieval-based synthetic tasks(Kamradt, 2023; Mohtashami & Jaggi, 2023; Wu et al., 2023; Li et al., 2023a; Liu et al., 2024a), with a few on other types of long-context usage, including various types of reasoning (Tay et al., 2021) and long-range discourse modeling (Sun et al., 2022).
Long-form Text Generation.
Significant advances have been made in long-form text generation, spanning various applications such as story generation (Fan et al., 2019b; Xu et al., 2020), paragraph completion (Kang & Hovy, 2020), sustained conversation (Xu et al., 2022), comprehensive question answering (Fan et al., 2019a; Dasigi et al., 2021; Stelmakh et al., 2022; Lee et al., 2023), and article production (Hua & Wang, 2020; Hu et al., 2022). Employing large language models (LLMs) for evaluative tasks (Liu et al., 2023; Chiang & Lee, 2023; Liu et al., 2024b; Bai et al., 2024) presents challenges, including limited access to current information and inconsistent performance (Shen et al., 2023). Evaluators often struggle to accurately assess quality in lengthy texts, leading to substantial costs (Gillick & Liu, 2010; Iskender et al., 2020). In contrast, the methodology proposed by Tan et al. (2024) aims to improve the assessment of informativeness and coverage, thus enhancing the evaluation of long-form text generation over extended narratives. However, the average length of text generated in the studies mentioned above does not reach the current context length limitations of long context language models, with most texts being fewer than 2000 tokens.
Instruction following.
Recent advancements in open-ended instruction tuning algorithms and models (Ouyang et al., 2022; Rafailov et al., 2024; OpenAI, 2022; Taori et al., 2023; Chiang et al., 2023) have underscored the need for scalable and cost-effective evaluation methods. Consequently, LLMs have been deployed as evaluators, showing better correlation with human assessments than traditional metrics such as BLEU (Papineni et al., 2002). Despite their advantages, LLMs exhibit biases including sensitivity to the order of presentation and a preference for verbose outputs from models with similar architectures (Wang et al., 2024; Pezeshkpour & Hruschka, 2023; Zheng et al., 2023). This has led to the development of meta-evaluation benchmarks like FairEval, MT-Bench, and LLMEval2(Wang et al., 2024; Zheng et al., 2023; Zhang et al., 2023) to address these biases. Recent studies have focused on improving LLM evaluations by employing diverse evaluators and enhanced prompting strategies(Zheng et al., 2023; Li et al., 2023c; Zhang et al., 2023; Chan et al., 2023). However, it is noteworthy that the length of instruction following in these studies often fails to meet the context length capabilities of current long-context language models.
6 Conclusion
We present LongGenBench , a synthetic benchmark designed to evaluate the long-form generation capabilities of language models. LongGenBench employs tasks that require following instructions over long distances, offering a comprehensive assessment of language models’ abilities. We evaluated nine advanced language models using LongGenBench, with context sizes ranging from 32K to 128K. Although these models achieved perfect scores in the commonly used ”Ruler”, their performance decreased significantly on LongGenBench. Common failure modes observed include premature termination of the main task, incomplete responses, disregard for prior instructions as outputs lengthened, and repetitive content generation. Our findings indicate that LongGenBench poses significant challenges for even the top-tier open-source models.
Appendix A Models
Our results in the main text only include aligned models (two closed-source model gpt-4o, gpt-4o-mini and eight open-source models, the model information is shown in the table 3.
Model Aligned Size Context Length Huggingface (Wolf et al., 2019) / API
GPT-4o-mini (OpenAI, 2024a) - 128K gpt-4-mini
GPT-4o (OpenAI, 2024b) - 128K gpt-4o-2024-08-06
Llama3.1-8B-Instruct (Dubey et al., 2024) 8B 128K meta-llama/Meta-Llama-3.1-8B-Instruct
Llama3.1-72B-Instruct (Dubey et al., 2024) 70B 128K meta-llama/Meta-Llama-3.1-70B-Instruct
Qwen2-7B-Instruct (Yang et al., 2024) 7B 128K Qwen/Qwen2-7B-Instruct
Qwen2-72B-Instruct (Yang et al., 2024) 72B 128K Qwen/Qwen2-72B-Instruct
Mistral-v0.2 (Mistral.AI, 2023) 7B 32K mistralai/Mistral-7B-Instruct-v0.2
Mixtral-8x7B (Jiang et al., 2024) 8x7B 32K mistralai/Mixtral-8x22B-Instruct-v0.1
FILM-7B (An et al., 2024b) 7B 128K In2Training/FILM-7B
Longwrite-llama3.1-8B (Bai et al., 2024) 8B 128K THUDM/LongWriter-llama3.1-8b
Table 3: Information of evaluated and analyzed models in LongGenBench.
Appendix B FOUR DISTINCT SCENARIO SETUPS
Here are four templates for different scenarios. In Diary template, the professions and names are interchangeable options.
Diary for 2018 Emma is a photographer with a passion for chronicling her vibrant life through weekly diary entries. Captures: 1) Single Instruction (SI): Birthdays of family members, wedding anniversaries, etc.; 2) Range Instruction (RI): Family beach vacation in Maui, Week-long road trip across the Pacific Coast Highway, etc.; 3) Periodic Instruction (PI): Attend golf lessons at the local club, Join a weekend hiking group, etc.; 4) Weekly updates on weather changes, work developments, family life, and other interesting topics. 5) Use ’#*#’ to separate each weekly entry (e.g. example) Generate a complete weekly diary for Emma for the entire year of 2018. Start from January 1st, a Monday, marking the first week, and continue through to December 31st, the end of the 52nd week. Ensure that the diary consists of 52 entries, one for each week. Each diary entry should be at least 200 words. When the design of all 52 weeks is complete, use ’*** finished ***’ to indicate the end of the document. Ensure clarity and continuity without any interruptions or omissions in the narrative throughout the year. *** started *** #*# Week 1 (January 1st - January 7th):
Menu for 2018 As Chef Roy, a world-renowned chef at a globally renowned restaurant, you are tasked with designing an entire year’s menu for 2018. Your menu should be varied and innovative, adhering to the following guidelines: 1) Single Instruction (SI): (”Independence Day Celebration”, ”2018-07-04”, ”American Apple Pie”), (”Summer Solstice Celebration”, None, ”Midsummer Night’s Fish Fry”), etc; 2) Range Instruction (RI): (”Mushroom Season Specials”, ”Various Mushroom Dishes”), (”Seafood Season Extravaganza”, ”Fresh Seafood Platter”), etc.; 3) Periodic Instruction (PI): (”Seafood Fridays”, 2, ”Fish and Chips”), (”Monthly Steak Night”, 3, ”Prime Ribeye Steak”), etc.; 4) Use ’#*#’ to separate each weekly menu (e.g. example) Generate a comprehensive weekly menu diary for the entire year of 2018, start from January 1st, a Monday, marking the first week, and continuing until December 31, the end of the 52nd week. Ensure that the diary consists of 52 entries, one for each week. Each weekly menu must include a detailed description of the offerings, featuring at least two options for appetizers, main courses, side dishes, desserts, and drinks. Ideally, between 200 and 220 words per menu description to ensure thoroughness and richness of detail. Conclude the diary with ’*** finished’ to signify the completion of the year’s menu planning. Ensure clarity and continuity without any interruptions or omissions in the menu throughout the year. *** started *** #*# Menu Week 1 (January 1st - January 7th):”,
Skyscraper Design Construct a skyscraper with 100 floors. Please follow the detailed floor assignments below: 1) Single Instruction (SI): office, conference room, retail store, etc; 2) Range Instruction (RI): hospital with various departments, corporate headquarters for a major company, etc.; 3) Periodic Instruction (PI): outdoor terrace, sky garden, etc.; 4) Document each floor independently with detailed descriptions of the intended facilities, architectural features, and unique design elements. 5) Use ’#*#’ to separate the documentation for each floor (e.g. example). Ensure that the document consists of 100 entries, each containing at least 150 words. Ensure clarity and continuity without any interruptions or omissions in the narrative throughout the document. When the design of all 100 floors is complete, use ’*** finished’ to indicate the end of the document. *** started *** #*# Floor 1:
Urban Planning Design a vibrant and diverse city using a 10x10 block grid, numbered from 1 to 100. Arrange the blocks sequentially from left to right and top to bottom. Ensure that each block is uniquely planned to reflect a wide array of city facilities, highlighting the rich urban environment and cultural diversity. 1) Single Instruction (SI): theater, museum, etc.; 2) Range Instruction (RI): shopping district, industrial park, etc.; 3) Periodic Instruction (PI): public restroom, convenience store, etc.; 4) Document each block independently with detailed descriptions of the intended facilities, architectural features, and unique design elements. 5) Use ’#*#’ to separate the documentation for each block like (e.g.example) Ensure that the document consists of 100 entries, each containing at least 150 words. Ensure that the document contains detailed descriptions for each block, with a minimum of 150 words per description. Ensure clarity and continuity in the narrative throughout the document without any interruptions or omissions. When all block assignments are complete, use ’*** finished’ to indicate the end of the document. *** started *** #*#Block 1 (0, 0):
Appendix C evaluation prompts
evaluation prompts Example 1 Example 2 Context: ”Corresponds to the output part of Specific Task Instruction” Instructions: Does this context include Specific Task Instruction? Please answer with ’yes’ or ’no’ only.”